# A Framework to Audit Email Address Privacy and Analyze Email Marketing Practices of Online Services and Apps

### v1

@@ Appendix B List of Domains
bloombergbusiness.com Financials 4 99 1 75 23
booking.com Consumer Discretionary 0 2 0 2 0
canva.com Information Technology 0 3 1 2 0
capcut.com Communication Services 0 69 19 49 1
cars.com Consumer Discretionary 0 15 6 9 0
cnet.com Information Technology 0 1 0 1 0
codepen.io Information Technology 0 49 1 47 1
codewars.com Information Technology 0 3 0 3 0
coursehero.com Consumer Discretionary 0 22 7 15 0
discogs.com Communication Services 0 8 0 7 1
discord.com Communication Services 0 3 0 3 0
doordash.com Consumer Discretionary 0 3 3 0 0
dunkinrewards.com Consumer Discretionary 0 70 61 8 1
ebay.com Consumer Discretionary 0 70 57 13 0
etsy.com Consumer Discretionary 4 196 115 81 0
expedia.com Consumer Discretionary 0 28 24 3 1
flickr.com Communication Services 0 44 13 31 0
foodnetwork.com Consumer Discretionary 0 1 0 1 0
foxsports.com Consumer Discretionary 4 111 39 63 9
fragrancenet.com Consumer Discretionary 0 50 41 9 0
freecodecamp.org Information Technology 0 1 0 1 0
geeksforgeeks.org Information Technology 0 12 7 5 0
github.com Information Technology 0 2 0 1 1
google.com Communication Services 0 3 0 2 1
hackerrankmail.com Information Technology 0 4 0 4 0
homedepot.com Consumer Discretionary 0 1 0 1 0
hotels.com Consumer Discretionary 0 89 72 17 0
imdb.com Consumer Discretionary 0 2 0 2 0
instagram.com Communication Services 0 14 0 13 1
kayak.com Consumer Discretionary 0 3 0 3 0
kbb.com Consumer Discretionary 0 1 0 1 0
khanacademy.org Information Technology 0 49 21 28 0
kohls.com Consumer Discretionary 3 335 330 5 0
leetcode.com Information Technology 0 8 5 3 0
linkedin.com Communication Services 0 31 1 30 0
lowes.com Consumer Discretionary 4 411 404 6 1
lyftmail.com Consumer Discretionary 0 7 1 6 0
microsoft.com Information Technology 0 3 0 3 0
nationalgeographic.com Consumer Discretionary 0 1 0 1 0
newegg.com Consumer Discretionary 0 1 1 0 0
nike.com Consumer Discretionary 0 1 0 1 0
openai.com Information Technology 0 3 0 3 0
opentable.com Consumer Discretionary 0 2 0 2 0
oracle-mail.com Information Technology 0 5 1 4 0
oracle.com Information Technology 0 1 0 1 0
orbitz.com Consumer Discretionary 0 15 14 1 0
pandora.com Communication Services 0 36 17 19 0
pinterest.com Communication Services 0 33 1 32 0
planetfitness.com Consumer Discretionary 0 1 0 1 0
plex.tv Consumer Discretionary 0 4 0 4 0
pons.com Consumer Discretionary 0 1 0 1 0
pornhub.com Consumer Discretionary 0 3 0 3 0
quizlet.com Consumer Discretionary 0 5 0 5 0
quora.com Information Technology 0 66 0 63 3
realtor.com Consumer Discretionary 0 7 2 5 0
reddit.com Communication Services 0 1 0 1 0
redditmail.com Communication Services 4 85 0 65 20
reuters.com Communication Services 0 2 0 2 0
ring.com Consumer Discretionary 0 5 1 4 0
roblox.com Consumer Discretionary 0 1 0 1 0
roku.com Communication Services 0 3 0 3 0
scrimba.com Information Technology 0 17 12 4 1
shazam.com Communication Services 0 1 0 1 0
shein.com Consumer Discretionary 0 106 93 12 1
sheinnotice.com Consumer Discretionary 0 1 0 1 0
shopify.com Consumer Discretionary 0 1 0 1 0
shutterstock.com Consumer Discretionary 0 52 20 31 1
skillcrush.com Consumer Discretionary 0 27 12 15 0
snapchat.com Communication Services 0 5 0 5 0
soundcloud.com Communication Services 0 11 3 8 0
southwest.com Consumer Discretionary 0 1 0 1 0
springernature.com Consumer Discretionary 0 2 0 2 0
statista.com Information Technology 0 1 0 1 0
steampowered.com Consumer Discretionary 0 1 0 1 0
target.com Consumer Discretionary 4 109 87 22 0
temu.com Consumer Discretionary 0 3 0 3 0
temuemail.com Consumer Discretionary 0 61 40 20 1
thumbtack.com Consumer Discretionary 0 11 1 10 0
tiktok.com Communication Services 0 1 0 1 0
today.com Information Technology 4 116 112 3 1
topcoder.com Information Technology 0 1 0 1 0
tracfone.com Consumer Discretionary 0 10 5 5 0
trip.com Consumer Discretionary 0 2 0 2 0
tripadvisor.com Consumer Discretionary 0 74 32 42 0
trivago.com Consumer Discretionary 0 32 26 6 0
turo.com Consumer Discretionary 0 1 0 1 0
tutorialsdojo.com Information Technology 0 21 15 5 1
tutorialspoint.com Information Technology 0 23 16 7 0
twitch.tv Communication Services 0 6 0 6 0
uber.com Consumer Discretionary 4 173 169 4 0
udacity.com Information Technology 0 1 0 1 0
usnews.com Consumer Discretionary 0 2 0 2 0
vimeo.com Communication Services 0 14 10 4 0
walmart.com Consumer Discretionary 0 1 1 0 0
wayfair.com Consumer Discretionary 1 549 520 27 2
waze.com Information Technology 0 2 0 2 0
webmd.com Health 4 270 9 250 11
wish.com Consumer Discretionary 1 501 455 46 0
xhamster.com Consumer Discretionary 0 1 0 1 0
xvideos.com Consumer Discretionary 0 1 0 0 1
yelp.com Consumer Discretionary 0 11 0 11 0
zillow.com Consumer Discretionary 0 1 0 1 0

### v2

@@ Appendix B List of Received Email Domains
bloombergbusiness.com Financials 0 99 1 75 23
booking.com Online Marketplace 0 2 0 2 0
canva.com Digital Services 0 3 1 2 0
capcut.com Communication Platforms 0 71 19 49 1
cars.com E-tailer 0 15 6 9 0
cnet.com Digital Services 0 1 0 1 0
codepen.io Digital Services 0 49 1 47 1
codewars.com Digital Services 0 3 0 3 0
coursehero.com Digital Services 0 22 7 15 0
discogs.com Communication Platforms 0 8 0 7 1
discord.com Communication Platforms 0 3 0 3 0
doordash.com Online Marketplace 0 3 3 0 0
dunkinrewards.com Brick and Mortar 0 70 61 8 1
ebay.com E-tailer 0 70 57 13 0
etsy.com E-tailer 1 196 115 81 0
expedia.com Digital Services 0 28 24 3 1
flickr.com Communication Platforms 0 44 13 31 0
foodnetwork.com Online Entertainment 0 1 0 1 0
foxsports.com Online Entertainment 0 111 39 63 9
fragrancenet.com E-tailer 0 50 41 9 0
freecodecamp.org Digital Services 0 1 0 1 0
geeksforgeeks.org Digital Services 0 12 7 5 0
github.com Digital Services 0 2 0 1 1
google.com Communication Platforms 0 3 0 2 1
hackerrankmail.com Digital Services 0 4 0 4 0
homedepot.com Omnichannel 0 1 0 1 0
hotels.com Online Marketplace 0 89 72 17 0
imdb.com Online Entertainment 0 2 0 2 0
instagram.com Communication Platforms 0 14 0 13 1
kayak.com Digital Services 0 3 0 3 0
khanacademy.org Digital Services 0 49 21 28 0
kohls.com Omnichannel 1 335 330 5 0
leetcode.com Digital Services 0 8 5 3 0
linkedin.com Communication Platforms 0 31 1 30 0
lowes.com Omnichannel 1 411 404 6 1
lyftmail.com Online Marketplace 0 7 1 6 0
nationalgeographic.com Online Entertainment 0 1 0 1 0
newegg.com E-tailer 0 1 1 0 0
nike.com Omnichannel 0 1 0 1 0
openai.com Digital Services 0 3 0 3 0
opentable.com Digital Services 0 2 0 2 0
oracle-mail.com Digital Services 0 5 1 4 0
oracle.com Digital Services 0 1 0 1 0
orbitz.com Online Marketplace 0 15 14 1 0
pandora.com Communication Platforms 0 36 17 19 0
pinterest.com Communication Platforms 0 33 1 32 0
planetfitness.com Brick and Mortar 0 1 0 1 0
plex.tv Online Entertainment 0 4 0 4 0
pons.com Digital Services 0 1 0 1 0
pornhub.com Online Entertainment 0 3 0 3 0
quizlet.com Digital Services 0 5 0 5 0
quora.com Digital Services 0 67 0 63 3
realtor.com Online Marketplace 0 7 2 5 0
reddit.com Communication Platforms 0 1 0 1 0
redditmail.com Communication Platforms 0 86 0 65 20
reuters.com Communication Platforms 0 2 0 2 0
ring.com Omnichannel 0 5 1 4 0
roblox.com Online Entertainment 0 1 0 1 0
roku.com Communication Platforms 0 3 0 3 0
scrimba.com Digital Services 0 17 12 4 1
shazam.com Communication Platforms 0 1 0 1 0
shein.com E-tailer 0 106 93 12 1
sheinnotice.com E-tailer 0 1 0 1 0
shopify.com Online Marketplace 0 1 0 1 0
shutterstock.com Digital Services 0 52 20 31 1
skillcrush.com Online Entertainment 0 27 12 15 0
snapchat.com Communication Platforms 0 5 0 5 0
soundcloud.com Communication Platforms 0 11 3 8 0
southwest.com Brick and Mortar 0 1 0 1 0
springernature.com Omnichannel 0 2 0 2 0
statista.com Digital Services 0 1 0 1 0
steampowered.com E-tailer 0 1 0 1 0
target.com Omnichannel 0 109 87 22 0
temu.com E-tailer 0 3 0 3 0
temuemail.com E-tailer 0 61 40 20 1
thumbtack.com Online Marketplace 0 11 1 10 0
tiktok.com Communication Platforms 0 1 0 1 0
today.com Digital Services 0 116 112 3 1
topcoder.com Digital Services 0 1 0 1 0
tracfone.com Brick and Mortar 0 10 5 5 0
trip.com Online Marketplace 0 2 0 2 0
tripadvisor.com Online Marketplace 0 74 32 42 0
trivago.com Online Marketplace 0 32 26 6 0
turo.com Online Marketplace 0 1 0 1 0
tutorialsdojo.com Digital Services 0 21 15 5 1
tutorialspoint.com Digital Services 0 23 16 7 0
twitch.tv Communication Platforms 0 6 0 6 0
uber.com Online Marketplace 0 174 169 4 0
udacity.com Digital Services 0 1 0 1 0
usnews.com Digital Services 0 2 0 2 0
vimeo.com Communication Platforms 0 14 10 4 0
walmart.com Omnichannel 0 1 1 0 0
wayfair.com E-tailer 1 559 520 27 2
waze.com Digital Services 0 2 0 2 0
webmd.com Digital Services 1 270 9 250 11
wish.com E-tailer 1 502 455 46 0
xhamster.com Online Entertainment 0 1 0 1 0
xvideos.com Online Entertainment 0 1 0 0 1
yelp.com Digital Services 0 11 0 11 0
zillow.com Online Marketplace 0 1 0 1 0